### v4

@@ 1 Introduction
Figure 1: Two common approaches to encode an example of four classical data features $X=x_{1},x_{2},x_{3},x_{4}$. In Figure 1(a), the four classical features are mapped into the rotation angles of the four one-qubit $R_{x}$ rotation gate (Angle Encoding). In Figure 1(b), the features are mapped into the amplitude of the two-qubits state |$\ket{\psi}$ (Amplitude Encoding). To prepare this quantum state, one-qubit $R_{y}$ rotation gates and CNOT gates must be applied [28], where the angles $\theta_{1},\theta_{2},\theta_{3}$ depend on the four classical data $X=x_{1},x_{2},x_{3},x_{4}$.
In this paper, we propose a hybrid Quantum Machine Learning model that integrates Parameterized Quantum Circuits (PQC) into a neural network framework. PQCs inherently preserve unitarity due to quantum mechanical principles, eliminating the need for computationally intensive orthogonality constraints. Moreover, we leverage amplitude encoding in PQCs for exponential data compression, reducing the number of required qubits logarithmically with input size. Compared to Angle Encoding, this approach minimizes trainable parameters, enhances efficiency, and maintains accuracy in high-dimensional settings.
The proposed method demonstrated superior performance using a global dataset of 1,725 observations with 256 features spanning 576 firms from 1996 to 2023. It achieved a Root Mean Square Error (RMSE) of 0.228, lower than the RMSE of 0.246 for classical Neural Networks and 0.242 for quantum models with Angle Encoding (see the Results section below for more details). Additionally, the QML model with Amplitude Encoding has fewer trainable parameters, leading to a faster training time of 0.73 seconds per epoch compared to the 0.81 necessary for the QML with Angle Encoding. The lower qubit requirements and reduced computation time underscore the practical applicability of our method for recovery rate forecasting.
The remainder of this paper is organized as follows: Section 2 presents the data and its features. Section 3 details the hybrid QML approach and amplitude encoding. Section 4 discusses the numerical analysis, and Section 5 concludes with insights and directions for future research.
2 Data
We consider a dataset comprising 256 features and 1,725 observations covering 576 firms from 1996 to 2023. The data is obtained through the NRF Research Project UP5 of the National University of Singapore. The UP5 data contains Macroeconomic and market-related features obtained from FRED and Refinitiv; financial statement features of firms sourced from Bloomberg (BBG); and bond-level features, as well as firm-level or market-level credit product features provided by the Credit Research Initiative (CRI) of the National University of Singapore.
[…]
$N$ Obs. Mean STD Min 25% 50% 75% Max
Recovery
Rate
1,725 0.4845 0.3317 0 0.1811 0.4170 0.7896 1.0996
Table 1: Summary Statistics of Recovery Rate
3 Methodology
Our Quantum Machine Learning (QML) model integrates a classical master layer and a Quantum Neural Network (QNN) to predict recovery rates, similar to the QML frameworks in [38]. The classical master layer comprises an input layer and a hidden layer of equal size, using a LeakyReLU activation function. This architecture offers two advantages. First, it extracts meaningful internal representations from high-dimensional, redundant input data. Second, it introduces non-linearity, enabling the model to capture complex relationships. Building on this classical layer, the QNN combines a Parameterized Quantum Circuit (PQC) and quantum data encoding to process the data.
3.1 Parameterized Quantum Circuit (PQC)
The PQC is a quantum circuit with adjustable parameters embedded in the rotation angles of single-qubit gates. To enhance expressiveness—the ability to represent diverse quantum states and span the Hilbert space [41]—entanglement between qubits is introduced using controlled gates, typically CNOT gates. A common PQC consists of:
@@ 3.1 Parameterized Quantum Circuit (PQC)
    A layer of $n$ single-qubit rotation gates, $G(\alpha,\beta,\gamma)$, applied to each qubit. The angles $\alpha,\beta,\gamma$ are the trainable parameters of the PQC.
@@
    A layer of $n$ two-qubits controlled gates to introduce entanglement. Specifically, we employ the CNOT gate with a range of one, where the $i$th qubit acts as the control and is connected to its adjacent $(i+1)$th qubit, which serves as the target.
This configuration, referred to as a strongly entangling circuit [39], provides the PQC with entangling power while maintaining a manageable number of parameters. Specifically, it requires $O(3n)$ trainable parameters for $n$ qubits. After the PQC computation, the expectation values of $Z$ Pauli observables are measured and passed to the classical output layer. A classical optimizer minimizes the loss function, Root Mean Squared Error (RMSE) in this case, and updates the network parameters via backpropagation.
3.2 Amplitude Encoding
Before the PQC processes input, classical data is encoded into quantum states, represented as |$\ket{\psi}$ (Figure 5). For the high-dimensional feature space in recovery rate prediction, we adopt Amplitude Data Encoding [28, 40, 39]. This technique maps $2^{n}$ classical features to the amplitudes of an $n$-qubit quantum state:
|$\ket{\psi}=\sum_{x\in\{0,1\}^{n}}\alpha_{x}\ket{x}$ (1)
where ${\alpha_{x}}$ are normalized amplitudes, calculated from the input data, satisfying:
$\sum_{x\in\{0,1\}^{n}}|\alpha_{x}|^{2}=1.$ (2)
@@ 3.2 Amplitude Encoding
Strongly Entangling PQC
Figure 5: The Strongly Entangling Layer PQC (Parameterized Quantum Circuit) begins with the quantum state input, denoted as |$\ket{\psi}$. Each rotation gate in the circuit is represented as $R(\alpha_{i},\beta_{i},\gamma_{i})$, where $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ are the rotation angles around the $X$, $Y$, and $Z$ axes of the Bloch sphere, respectively. These angles are the trainable parameters of the PQC.
A key advantage of Amplitude Encoding is its exponential compression of input data. The number of qubits required grows only logarithmically with the number of features, resulting in fewer trainable parameters in the PQC, improving the model’s scalability and efficiency. For instance, the number of trainable parameters in the strongly entangling PQC used in this study scales as $O(3\log_{2}N)$, making it well-suited for handling high-dimensional data.
Though some limitations due to noise and decoherence might affect the computation in real quantum hardware, in this work, we train the QML model in a fault-tolerant quantum simulator where decoherence or gate errors are not concerns. The complete PQC structure and data encoding process are illustrated in Figures 5 and 6.
[…]
PQCOutput
3.3 Alternative Quantum and Classic Model
To comprehensively evaluate the performance of our proposed QML model, we compare it against two alternative models: a classical feedforward neural network (FNN) and another QML model utilizing Angle Encoding, inspired by [38].
Angle Encoding [40, 37, 35, 15] maps classical data onto the rotation angles of single-qubit gates. This method requires one qubit per input feature, making it computationally demanding and impractical for high-dimensional datasets like those used in our recovery rate prediction due to the limited availability of logical qubits in current quantum hardware and classical simulator. To address this challenge, we follow the approach outlined in [38], incorporating a classical preprocessing layer (called auxiliary layer hereafter) to reduce the dimensionality of the input data. This auxiliary layer extracts key features and ensures that the number of features aligns with the number of available qubits. The outputs of this classical layer are then encoded into the quantum circuit using Angle Encoding. This architecture, illustrated in Figure 7, enables scalability by allowing flexibility in selecting the number of qubits while maintaining the model’s expressiveness. The number of trainable parameters in the additional auxiliary layer and the PQC scales as $O(3Nn)$, where $N$ is the number of input features, and $n$ is the number of selected qubits.
[…]
PQCOutput
Figure 7: The QML model with Angle Encoding. Starting from a set of $N$ input features, we introduce an auxiliary layer to reduce the number of inputs as the amount of $n$ qubits. The outputs of the auxiliary layer are classical data encoded into the angle of the single qubit rotation gate $R_{x}(x_{i})$. After the application of the Strongly Entangling PQC with Angle Encoding, a measurement is performed. These measurement results are then sent to the classical output layer and post-processed in the classical optimizer.
The classical feedforward neural network (FNN) serves as another benchmark for comparison. In the FNN, a hidden layer is appended to the master input layer, with the number of hidden nodes carefully chosen to ensure a comparable number of trainable parameters to the QML models. This setup not only provides a fair basis for comparison but also allows us to assess the effectiveness of the classical model and its susceptibility to overfitting. By tuning the number of hidden nodes, we balance the trade-off between model complexity and predictive performance.
Through these comparisons, we aim to evaluate the strengths and limitations of our QML model relative to classical neural networks and alternative quantum approaches, particularly in handling the intricate, high-dimensional relationships present in recovery rate prediction.
@@ 4 Results
In this section, we present the results produced by our models, based on the methodologies and parameter settings outlined in Sections 3 and 2. These findings offer insights into the predictive performance of both classical and quantum machine learning approaches on the chosen dataset. The evaluation of the models is conducted using standard metrics, such as RMSE calculated through k-fold cross-validation.
@@ 4.1 Parameter Settings and Experimental Setup
In our proposed QML architecture, the number of qubits is fixed, with 256 classical features encoded into 8 qubits. While additional qubits could theoretically be utilized with Amplitude Encoding, doing so introduces added complexities and challenges. These include an increase in trainable parameters, the need to handle padding during initialization to accommodate the extra qubits, deeper circuits for state preparation, and potential redundancy in the input data. In this study, since the simplest configuration with 256 features encoded in 8 qubits yielded satisfactory results, we opted to focus on this straightforward setup, leaving the exploration of more complex configurations for future research.
The parameters used in the regression models are optimized to minimize RMSE on the training data selected in a cross-validation setting. The optimization process utilized the Adam optimizer [19], with specific hyperparameters such as learning rate and batch size detailed in Table 2. These parameters were selected through a systematic grid search to ensure optimal model performance. Table 2 also lists the computational resources used for training and evaluation, including both classical and quantum setups.
@@
The quantum machine learning (QML) models were implemented using PennyLane (version 0.38.0), an open-source framework for quantum programming and differentiable PQCs. All QML models with PQCs were executed on a state-vector quantum simulator. Specifically, the built-in PennyLane device called default.qubit [31, 30]. The default.qubit device, written in Python with Autograd and PyTorch backends, simulates quantum operations and performs measurements on quantum systems using a classical CPU. In particular, we conducted our quantum experiments on the 13th Gen Intel® Core™ i9-13900H CPU.
@@
Table 2 summarizes the devices used for training the classical and quantum layers together with the hyperparameters of the Adam optimizer.
@@
Layer Device Gradient Optimizer Learning Rate Batch size
[…]
Table 2: Specification of the devices, the gradient calculation methods, the optimizer, and the hyperparameters for the classical and quantum layer.
4.2 Model Comparison and Evaluation
In this section, we evaluate the effectiveness of our proposed QML model using Amplitude Encoding (QML Amp), comparing it against a QML model with Angle Encoding (QML Ang) based on [38] and a classical feedforward neural network (FNN). The specifications of the models’ hyperparameter and architecture are reported in Table 3.
@@ 4.2 Model Comparison and Evaluation
Model Input Layer I Hidden Layer Auxiliary Layer II Hidden Layer
FNN 256 256 LeakyRelu with slope = -0.3 No 8 LeakyRelu with slope = -0.3
@@
QML Ang 256 256 LeakyRelu with slope = -0.3 8 8 Qubits Strongly Entangling PQC
@@
QML Amp 256 256 LeakyRelu with slope = -0.3 No 8 Qubits Strongly Entangling PQC
Table 3: The specification of the model’s architecture.
To benchmark the proposed models, we conduct k-fold cross-validation by dividing the dataset into four folds, each consisting of 75% training data and 25% test data. To ensure a fair comparison, all models and experiments are executed on the same four-fold splits for a fixed number of one hundred epochs. During training, we record the RMSE calculated on the test data of each fold. To achieve comparable trainable parameter counts across models, the number of qubits in the QML Ang the number of hidden nodes in the classical FNN are selected to approximate the parameter count of the QML Amp. A detailed discussion of how the performance of the FNN and the QML Ang evolves as the number of hidden nodes (for the FNN) and qubits (for the QML model) increases is provided in Appendices A and B.
Figure 8 presents the average RMSE (solid line) and the standard deviation of RMSE values (shaded area) achieved by the proposed models. Table 4 reports the best average RMSE on test data achieved during training, the average RMSE standard deviation, alongside the average execution time per epoch.
Figure 8: Average test RMSE and standard deviation (shaded area) calculated over four different cross-validation folds for the FNN, QML Ang, and QML Amp.
Model ##\## Parameters Best Average RMSE At epoch # Average STD Average Time per Epoch (s)
FNN 67,857 0.246 90 0.018 0.03
QML Ang 67,873 0.242 78 0.009 0.81
QML Amp 65,825 0.228 55 $\mathbf{0.008}$ 0.73

### v5

@@ 4.2 Quantum Data Encoding
Figure 5: Two common approaches to encode an example of four classical data features $X=x_{1},x_{2},x_{3},x_{4}$. In Figure 5(a), the four classical features are mapped into the rotation angles of the four one-qubit $R_{x}$ rotation gate (Angle Encoding). In Figure 5(b), the features are mapped into the amplitude of the two-qubits state |$\ket{\psi}$ (Amplitude Encoding). To prepare this quantum state, one-qubit $R_{y}(\bm{\theta})$ rotation gates and CNOT gates must be applied, where the angles $\bm{\theta}=\theta_{1},\theta_{2},\theta_{3}$ depend on the four classical data $X=x_{1},x_{2},x_{3},x_{4}$ (see [66] for furhter details).
The preparation of the state |$\ket{\psi}$ in Amplitude Encoding follows the protocol described in [67]. This method uses a sequence of single-qubit $R_{y}(\bm{\theta})$ rotations and CNOT gates to transform the initial zero state |$\ket{\bm{0}}$ into the desired amplitude-encoded state |$\ket{\psi}$. As illustrated in Figure 5, for a four-dimensional input vector $X=\{x_{1},x_{2},x_{3},x_{4}\}$, Angle Encoding requires four qubits (one per feature), while Amplitude Encoding only requires two qubits. This exponential compression enables us to scale to higher-dimensional datasets using fewer quantum resources. In our case, 256 input features are encoded using only eight qubits. This figure essentially depicts the interface between the classical layer (the pink block in Figure 4) and the quantum neural network (the purple block in Figure 4).
4.3 Parametrize Quantum Circuit (PQC)
Once encoded, the quantum state |$\ket{\psi}$ is passed through a Parameterized Quantum Circuit (PQC) in the solid block of the QNN in Figure 4. The PQC consists of a sequence of parameterized single-qubit rotation gates and, to increase expressibility—the capacity to represent diverse quantum states and explore the Hilbert space [89]—entangling CNOT gates. A typical PQC architecture includes:
@@ 4.3 Parametrize Quantum Circuit (PQC)
    A layer of $n$ single-qubit rotation gates, $R(\alpha,\beta,\gamma)$, applied to each qubit. Here, $\alpha,\beta,\gamma$ are the trainable parameters.
@@
    A layer of $n$ two-qubit entangling gates. Specifically, we use CNOT gates with a range of one: the $i$th qubit acts as the control for the $(i+1)$th target qubit.
This configuration, commonly referred to as a Strongly Entangling PQC [87], provides the PQC with entangling power while maintaining a manageable number of parameters, specifically $O(3n)$ trainable parameters for $n$ qubits. The overall architecture results in a smaller number of trainable parameters in the QNN, specifically scaling as $O(3\log_{2}N)$, where $N$ represents the number of features.
After the quantum computation, the expectation values of $Z$ Pauli observables are measured and passed to the classical output layer. The resulting measurement outcomes are then passed back to the classical domain, where they are post-processed by a final classical output layer for regression (i.e., recovery rate prediction).
This hybrid setup effectively leverages the strengths of both classical and quantum components: the classical neural layer captures coarse patterns and performs feature conditioning, while the QNN with Amplitude Encoding provides a compact and expressive mechanism for modeling nonlinear dependencies within a constrained qubit budget.
4.4 Alternative regression models for comparison evaluation
To thoroughly evaluate the performance of our proposed QML model, we compare it against both classical and quantum alternatives. The first natural baseline is a QML model using Angle Encoding, inspired by [84], which is briefly summarized in the following Section 4.4.1. Among classical models, the most direct counterpart for comparison is a Feedforward Neural Network (FNN), as our QML architecture closely resembles an FNN with a QNN replacing the hidden layer. In our comparative analysis, we also include a baseline XGBoost [21] regressor. Although this model does not mirror the architecture of our QML model, we include it due to its widespread adoption and strong empirical performance.
Table 4 summarizes the models considered in this study, along with the subsequent sections where the comparison analysis is presented.
Model Section
QML Angle Encoding–QML Amplitude Encoding  5.3
FNN–QML Amplitude Encoding  5.3
XGBoost–QML Amplitude Encoding  5.4
Table 4: The list of regression models considered for the comparison evaluation.
4.4.1 Quantum Machine Learning model with Angle Encoding
Angle Encoding [88, 83, 80, 39] maps classical data onto the rotation angles of single-qubit gates. This method requires one qubit per input feature, making it computationally demanding and impractical for high-dimensional datasets like those used in our recovery rate prediction due to the limited availability of logical qubits in current quantum hardware and classical simulator. To address this challenge, we follow the approach outlined in [84], incorporating a classical preprocessing layer (referred to as the auxiliary layer hereafter) to reduce the dimensionality of the input data. This auxiliary layer extracts key features and ensures that the number of variables aligns with the number of available qubits. The outputs of this classical layer are then encoded into the quantum circuit using Angle Encoding as in Figure 5(a). This architecture, illustrated in Figure 6, enables scalability by allowing flexibility in selecting the number of qubits while maintaining the model’s expressiveness. The number of trainable parameters in the additional auxiliary layer and the QNN scales as $O(3Nn)$, where $N$ is the number of input features, and $n$ is the number of selected qubits.
@@ 4.4.1 Quantum Machine Learning model with Angle Encoding
QNN Angle EncodingOutput
Figure 6: The QML model with Angle Encoding. Starting from a set of $N$ input features, we introduce an auxiliary (classical) layer that reduces the number of inputs to match the number of qubits $n$ used in the quantum circuit. For visualization purposes, the figure shows an example with three qubits, but the number of qubits can be adjusted as needed. The outputs of the auxiliary layer are classical values, which are encoded into quantum states using Angle Encoding via single-qubit rotation gates of the form $R_{x}(x_{i})$. After applying a Strongly Entangling Parameterized Quantum Circuit (PQC), a measurement is performed. The resulting measurement outcomes are then passed to a classical output layer and further processed by a classical optimizer.
@@ 5 Results
In this section, we present the results produced by our models, based on the methodologies and settings outlined in Sections 4 and Table 4. These findings offer insights into the predictive performance of both classical and QML approaches on the chosen dataset. The evaluation of the models is conducted using standard metrics, such as RMSE calculated through four-fold cross-validation, leave-one-out cross-validation, along with statistical significance tests to determine which model performs significantly better than the others.
@@ 5.1 Parameter Settings and Experimental Setup
The parameters in the QML and FNN regression models are optimized to minimize RMSE on the training data selected in a cross-validation setting. The optimization process utilized the Adam optimizer [51], with specific hyperparameters such as learning rate and batch size detailed in Table 5. These hyperparameters were selected through a systematic grid search to ensure optimal model performance. Table 5 also lists the computational resources used for training and evaluation, including both classical and quantum setups.
@@
The QML models were implemented using PennyLane (version 0.38.0), an open-source framework for quantum programming and differentiable PQCs. All QML models with PQCs were executed on a state-vector quantum simulator. Specifically, the built-in PennyLane device called default.qubit [74, 73]. The default.qubit device, written in Python with Autograd and PyTorch backends, simulates quantum operations and performs measurements on quantum systems using a classical CPU. In particular, we conducted our quantum experiments on the 13th Gen Intel® Core™ i9-13900H CPU.
@@
In addition to the FNN and QML models, XGBoost regressor is also included for its robustness and strong predictive performance on tabular data. We enabled GPU acceleration using the XGBoost Python library (version 3.0.0). For XGBoost, both the L2 regularization strength and the learning rate were tuned via grid search.
Table 5 summarizes the hardware used for model training, optimization procedures, and key hyperparameters.
@@
Model Device Module Optimization Gradient Batch Size Learning Rate
XGBoost GPU XGBoost Newton-Raphson None None 0.1
FNN GPU PyTorch Adam Backpropagation 64 0.001
QML default.qubit state vector simulator on CPU Pennylane (PyTorch backend) Adam Backpropagation 64 0.001
Table 5: Summary of device specifications, optimization protocols, and hyperparameters for each regression model.
5.2 Models’ Architecture and Benchmark Metrics
In our proposed QML architecture, the number of qubits is fixed, with 256 classical features encoded into eight qubits using Amplitude Encoding. While it is theoretically possible to use more qubits, doing so introduces additional complexities, such as an increased number of trainable parameters, the need for input padding during initialization, deeper circuits for state preparation, and potential redundancy in the input representation. Since the simplest configuration—encoding 256 features into eight qubits—yielded satisfactory results, we chose to focus on this setup, leaving the exploration of more complex architectures for future work.
For the alternative models, particularly the FNN and the QML with Angle Encoding, the number of hidden nodes and qubits can be tuned to improve performance. However, to ensure a fair comparison, we selected the number of hidden nodes in the FNN and the number of qubits in the QML with Angle Encoding to match the trainable parameter count of the QML with Amplitude Encoding. Notably, varying these values did not significantly impact performance. A detailed analysis of how performance evolves with the number of hidden nodes in the FNN and qubits in the QML with Angle Encoding is provided in Appendix A.1 and Appendix A.2 respectively.
For the XGBoost regression model, tree depth was tuned via grid search, with a depth of six yielding the best results. All models were trained for a fixed number of 100 epochs, which for XGBoost corresponds to sequential updates of 100 decision trees. Further performance analysis of XGBoost with varying tree depth and number is presented in Appendix A.3. A summary of the model architectures and hyperparameters is provided in Table 6.
@@ 5.2 Models’ Architecture and Benchmark Metrics
Model Input Layer I Hidden Layer Auxiliary Layer II Hidden Layer Training Epochs / Trees
XGBoost 256 None None None 100 Boosted Trees (depth = 6)
FNN 256 256 LeakyReLU (slope = -0.3) None 8 LeakyReLU (slope = -0.3) 100 Epochs
@@
QML Angle Encoding 256 256 LeakyReLU (slope = -0.3) 8 8 Qubits Strongly Entangling PQC 100 Epochs
@@
QML Amplitude Encoding 256 256 LeakyReLU (slope = -0.3) None 8 Qubits Strongly Entangling PQC 100 Epochs
Table 6: Specifications of model architectures, including training epochs or number of trees for each model.
To benchmark the proposed models, we perform four-fold cross-validation, dividing the dataset into four folds, each comprising 75% training data and 25% test data. To ensure a fair comparison, all models are evaluated using the same four-fold data splits. During training, we record the root mean squared error (RMSE) on the test set for each fold and epoch to monitor the convergence behavior of each model.
Statistical significance testing is essential to ensure that observed performance differences between models reflect meaningful and robust improvements rather than random variation, as emphasized in recent studies such as [54, 43], and through established methods like the Diebold-Mariano test [28]. To assess whether performance differences between models are statistically significant, we conduct a Diebold-Mariano (DM) test, under the assumption of independent forecast errors across observations. For each pair of models (hereafter indicated as $a$ and $b$), we compute the $DM_{a,b}$ statistic across all folds and epochs as follows
$DM_{a,b}=\frac{\bar{d}_{a,b}}{\sqrt{\mathrm{Var}(\bar{d}_{a,b})}},$ (2)
where $\bar{d}_{a,b}$ indicates the mean of the differences between the absolute value of the residuals of the model $a$ and $b$, and $\mathrm{Var}(\bar{d}_{a,b})$ is the estimated variance of the mean difference.
The DM statistic allows us to test whether model $a$ is significantly better than model $b$. A negative value of $DM_{a,b}$ indicates that model $a$ performs better than model $b$, and vice versa for a positive value. Given the large number of test observations (over 400), we compare the DM statistic to the critical values 0.05 from a standard normal distribution to assess the statistical significance of the null hypothesis $DM_{a,b}=0$, and then conclude that the performance between model $a$ and $b$ is significantly different.
To further strengthen our evaluation, we also perform leave-one-out cross-validation (LOOCV) as an additional robustness check. As with four-fold cross-validation, RMSE is tracked at each epoch for every left-out observation to monitor convergence and model performance.
5.3 Comparison evaluation with Neural Netwoks
5.3.1 Cross and Leave-One-Out-Cross Validation
Figures 7(a) and 7(b) show the average RMSE (solid lines) and the corresponding standard deviations (shaded areas) obtained by the FNN, QML with Angle Encoding, and QML with Amplitude Encoding models under four-fold and leave-one-out cross-validation, respectively. Table 7 reports the best average test RMSE achieved during training, the associated average RMSE standard deviation, and the average execution time per epoch.
(a) Four-fold cross validation (CV)
(b) Leave-one-out cross validation (LOOCV)
Figure 7: Average test RMSE (solid lines) and standard deviation (shaded area) calculated on a four-fold cross validation (CV left panel) and leave-one-out cross validation (LOOCV right panel) for the FNN, QML with Angle Encoding, and QML with Amplitude Encoding. For the sake of visualization, given the high variance of the RMSE calculated on each left-out observation, the shaded area on the right panel indicates a scaled variance by a factor of four.
Model # Parameters Avg. Time per Epoch (s) Best Avg. RMSE At Epoch # Avg. STD
FNN 67,857 0.03 0.246 (CV) 90 (CV) 0.018 (CV)
0.160 (LOOCV) 93 (LOOCV) 0.044 (LOOCV)
QML Angle 67,873 0.81 0.242 (CV) 78 (CV) 0.009 (CV)
Encoding 0.153 (LOOCV) 90 (LOOCV) 0.044 (LOOCV)
QML Amplitude 65,825 0.73 0.228 (CV) 55 (CV) 0.008 (LOOCV)
Encoding 0.136 (LOOCV) 86 (LOOCV) 0.043 (LOOCV)